# Certifying zeros of polynomial systems using interval arithmetic

### v5

@@ 2.2 Complex interval arithmetic
where $X+iY=\{x+iy\mid x\in X,y\in Y\}$ and $i=\sqrt{-1}$. Following [May17, Ch. 9] we define the algebraic operations for $I=X+iY,J=W+iZ\in\mathbb{IC}$ in terms of operations on the real intervals from (1):
@@ 3.1 Krawczyk’s method
    If additionally $\lVert\mathbf{1}_{n}-Y\square\mathrm{J}F(I)\rVert_{\infty}<1$, then $F$ has exactly one zero in $I$.
@@
Let $F:\mathbb{C}^{n}\rightarrow\mathbb{C}^{n}$ be differentiable and $I\in\mathbb{IC}^{n}$. Let $K_{x,Y}(I)$ be the associated Krawczyk operator (see Definition 3.3). If there exists an invertible matrix $Y\in\mathbb{C}^{n\times n}$, such that $K_{x,Y}(I)\subset I$, we say that $I$ is an interval approximate zero $F$. We call $I$ a strong interval approximate zero of $F$ if in addition $\lVert\mathbf{1}_{n}-Y\square\mathrm{J}F(I)\rVert_{\infty}<1$
@@ 4.3 Determining strong interval approximate zeros
In a first step, the certification routine attempts to produce for a given $x\in X$ a strong interval approximate zero $I\in\mathbb{IC}^{n}$. Recall that for $I\in\mathbb{IC}^{n}$ to be a strong interval approximate zero we need by Theorem 3.6 to have a point $\tilde{x}\in I$, and a matrix $Y\in\mathbb{C}^{n\times n}$ such that $K_{\tilde{x},Y}(I)\subset I$, and $\lVert\mathbf{1}_{n}-Y\square\mathrm{J}F(I)\rVert_{\infty}<1$.
@@ 6 Proof of the Krawczyk method
Before we give the proof of Theorem 3.6, we need a lemma. It is a direct sequence of a complex version of the mean-value theorem.
@@
    $G_{Y}(z)-G_{Y}(x)\in(\mathbf{1}_{n}-Y\cdot\square\mathrm{J}F(I))\,(z-x).$
@@
In the proof we abbreviate $G:=G_{Y}$. We first show the second part assuming the first part of the lemma. Then, we prove the first part. We fix an interval $I\in\mathbb{IC}^{n}$ and $x,z\in\mathbb{C}^{n}$.
For the second part, we have to show that for all $I\in\mathbb{IC}^{n}$ we have $G(I)\subset K_{x,Y}(I)$. To show this we define the interval matrix $M:=(\mathbf{1}_{n}-Y\square\mathrm{J}F(I))\in\mathbb{IC}^{n\times n}$. The definition of $K_{x,Y}$ implies $G(x)+M(I-x)\subseteq K_{x,Y}(I)$. The first part of the lemma shows that for all $z\in I$ we have $G(z)-G(x)\in M(z-x)$. Consequently, $G(z)\in G(x)+M(I-x)\subseteq K_{x,Y}(I)$. This finishes the proof for the second part.
@@
The first part of the lemma may be shown entry-wise. We will show this by combining the mean value theorem with the following observation: $\mathrm{J}G(x)=\mathbf{1}_{n}-Y\cdot\mathrm{J}F(x)$, so we have the inclusion
$\mathrm{J}G(I)=\mathbf{1}_{n}-Y\cdot\mathrm{J}F(I)\subseteq\mathbf{1}_{n}-Y\cdot\square\mathrm{J}F(I).$ (4)
We relate $G(z)-G(x)$ to (4) using the mean value theorem. For this, we fix $1\leq j\leq n$. Let us first assume that $z=(z_{1},\ldots,z_{n})=(1,0,\ldots,0)$ and $x=(x_{1},\ldots,x_{n})=(0,0\ldots,0)$. In this case, the complex mean value theorem implies that there are $c_{j,1},c_{j,2}$ on the line segment between $1$ and 00, such that
$\mathrm{Re}(\tfrac{\mathrm{d}}{\mathrm{d}x_{1}}G_{j}(c_{j,1}))=\mathrm{Re}(G_{j}(z)-G_{j}(x)),\quad\mathrm{Im}(\tfrac{\mathrm{d}}{\mathrm{d}x_{1}}G_{j}(c_{j,2}))=\mathrm{Im}(G_{j}(z)-G_{j}(x)).$
This yields
$G_{j}(z)-G_{j}(x)$ $=(\mathrm{Re}(\tfrac{\mathrm{d}}{\mathrm{d}x_{1}}G_{j}(c_{j,1}))+i\cdot\mathrm{Im}(\tfrac{\mathrm{d}}{\mathrm{d}x_{1}}G_{j}(c_{j,2})))\,(z_{1}-x_{1})$
$=(\mathrm{Re}(G_{j}^{\prime}(c_{j,1}))+i\cdot\mathrm{Im}(G_{j}^{\prime}(c_{j,2})))\,(z-x),$
where $G_{j}^{\prime}(x)\in\mathbb{C}^{1\times n}$ denotes the complex derivative of $G_{j}$ at $x\in\mathbb{C}^{n}$. Putting these equations ranging over $j$ together we get
$G(z)-G(x)=A(z-x),$
where the $j$-th row of $A$ is $\mathrm{Re}(G_{j}^{\prime}(c_{j,1})+i\cdot\mathrm{Im}(G_{j}^{\prime}(c_{j,2})).$ By construction, $c_{j,1}$ and $c_{j,2}$ are contained in $I$, because $x$ and $z$ are contained in $I$, and $I$ is convex. Combined with (4) this yields the desired equality $G(z)-G(x)\in(\mathbf{1}_{n}-Y\cdot\square\mathrm{J}F(I))(z-x).$
For general $z$ and $x$ we find an invertible affine-linear transformation $\varphi:\mathbb{C}^{n}\to\mathbb{C}^{n}$ such that $\varphi(1,0,\ldots,0)=z$ and $\varphi(0,0,\ldots,0)=x$. Notice that $\varphi^{-1}(I)$ is convex, since $I$ is convex, and that convexity was the only property of $I$ that we have used in the previous paragraph. Therefore, we can apply what we have shown above to $H:=G\circ\varphi$ and get
$G(z)-G(x)\in(\mathbf{1}_{n}-Y\cdot\square\mathrm{J}H(\varphi^{-1}(I)))(\varphi^{-1}(z-x))=(\mathbf{1}_{n}-Y\cdot\square\mathrm{J}G(I))(z-x),$
as was stated. ∎
@@
$G_{Y}(z_{0})-G_{Y}(z_{1})\in(\mathbf{1}_{n}-Y\cdot\square\mathrm{J}F(I))\,(z_{0}-z_{1}).$
Applying norms and using submultiplicativity yields
$\lVert G_{Y}(z_{0})-G_{Y}(z_{1})\rVert_{\infty}\leq\lVert(\mathbf{1}_{n}-Y\cdot\square\mathrm{J}F(I))\rVert_{\infty}\,\lVert z_{0}-z_{1}\rVert_{\infty}.$ (5)
By assumption $\lVert\mathbf{1}_{n}-Y\cdot\square\mathrm{J}F(I)\rVert_{\infty}$ is smaller than 1 so $G_{Y}$ is a contraction. Banach’s Fixed Point Theorem implies that $G_{Y}$ has a unique fixed point in $I$. Since $Y$ is invertible, this fixed point is the unique zero of $F$ in $I$. This shows the second part of Theorem 3.6.
Finally, let $(x_{i})_{i\geq 0}$ be the sequence defined in Proposition 3.10. By assumption, $x_{0}\in I$ and $x_{i+1}=G_{Y}(x_{i})$, $i\geq 0$. The second part of Lemma 6.1 together with an induction argument imply that $x_{i}\in I$ for all $i$. Let $x^{*}$ be the unique zero of $F$ in $I$. Then, 5 holds if we replace $z_{0}$ by $x^{*}$ and $z_{1}$ by $x_{i}$. We get $\lVert x^{*}-x_{i+1}\rVert_{\infty}<c\cdot\lVert x^{*}-x_{i}\rVert_{\infty}$ with a constant $c:=\lVert\mathbf{1}_{n}-Y\cdot\square\mathrm{J}F(I)\rVert_{\infty}<1$. This shows linear convergence and finishes the proof of Proposition 3.10. ∎
@@ 7 Certifying reality
Let $F:\mathbb{C}^{n}\rightarrow\mathbb{C}^{n}$ be a real square system of polynomials (or rational functions) and $I\in\mathbb{IC}^{n}$ a strong interval approximate zero of $F$. Then there exists $x\in I$ and $Y\in\mathbb{C}^{n\times n}$ satisfying $K_{x,Y}(I)\subset I$ and $\lVert\mathbf{1}_{n}-Y\square\mathrm{J}F(I)\rVert_{\infty}<1$. If additionally $\{\bar{z}\,|\,z\in K_{x,Y}(I)\}\subset I$, the associated zero of $I$ is real.

### v6

@@ 2.2 Complex interval arithmetic
where $X+iY=\{x+iy\mid x\in X,y\in Y\}$ and $i=\sqrt{-1}$. In particular, every rectangular interval $I\in\mathbb{IC}$ has the property that
$x+iy,a+ib\in I\quad\Longrightarrow\quad x+ib\in I.$ (2)
Following [May17, Ch. 9] we define the algebraic operations for $I=X+iY,J=W+iZ\in\mathbb{IC}$ in terms of operations on the real intervals from (1):
@@ 3.1 Krawczyk’s method
    If additionally $\sqrt{2}\,\lVert\mathbf{1}_{n}-Y\square\mathrm{J}F(I)\rVert_{\infty}<1$, then $F$ has exactly one zero in $I$.
@@
Let $F:\mathbb{C}^{n}\rightarrow\mathbb{C}^{n}$ be differentiable and $I\in\mathbb{IC}^{n}$. Let $K_{x,Y}(I)$ be the associated Krawczyk operator (see Definition 3.3). If there exists an invertible matrix $Y\in\mathbb{C}^{n\times n}$, such that $K_{x,Y}(I)\subset I$, we say that $I$ is an interval approximate zero $F$. We call $I$ a strong interval approximate zero of $F$ if in addition $\sqrt{2}\lVert\mathbf{1}_{n}-Y\square\mathrm{J}F(I)\rVert_{\infty}<1$
@@ 4.3 Determining strong interval approximate zeros
In a first step, the certification routine attempts to produce for a given $x\in X$ a strong interval approximate zero $I\in\mathbb{IC}^{n}$. Recall that for $I\in\mathbb{IC}^{n}$ to be a strong interval approximate zero we need by Theorem 3.6 to have a point $\tilde{x}\in I$, and a matrix $Y\in\mathbb{C}^{n\times n}$ such that $K_{\tilde{x},Y}(I)\subset I$, and $\sqrt{2}\,\lVert\mathbf{1}_{n}-Y\square\mathrm{J}F(I)\rVert_{\infty}<1$.
@@ 6 Proof of the Krawczyk method
Before we give the proof of Theorem 3.6, we need a lemma. It is a consequence of a complex version of the mean-value theorem.
@@
    $\|G_{Y}(z)-G_{Y}(x)\|_{\infty}\leq\sqrt{2}\,\|\mathbf{1}_{n}-Y\cdot\square\mathrm{J}F(I)\|_{\infty}\,\|z-x\|_{\infty}.$
@@
In the proof we abbreviate $G:=G_{Y}$. We fix an interval $I\in\mathbb{IC}^{n}$ and $x,z\in\mathbb{C}^{n}$.
@@
The first part of the lemma may be shown entry-wise. We will show this by combining a variant of the mean value theorem with the following observation: $\mathrm{J}G(x)=\mathbf{1}_{n}-Y\cdot\mathrm{J}F(x)$, so we get, using 2, the inclusion
$\mathrm{Re}(\mathrm{J}G(I))+i\,\mathrm{Im}(\mathrm{J}G(I))\subseteq\mathbf{1}_{n}-Y\cdot\square\mathrm{J}F(I).$ (5)
We define $w:=\mathrm{Re}(z)+i\mathrm{Im}(x)$. Let $1\leq j\leq n$ and let $G_{j}$ denote the $j$-th entry of $G$. We define the function $h(t):=G_{j}(tz+(1-t)w)$. The real and imaginary part of $h(t)$ are real differentiable functions of the real variable $t$. The mean value theorem can be applied, and we find $10<t_{1},t_{2}<$ 1 such that $\mathrm{Re}(h(1))-\mathrm{Re}(h(0))=\tfrac{\mathrm{d}}{\mathrm{d}t}\mathrm{Re}(h(t_{1}))$ and $\mathrm{Im}(h(1))-\mathrm{Im}(h(0))=\frac{\mathrm{d}}{\mathrm{d}t}\mathrm{Im}(h(t_{2}))$. Setting $c_{1}=t_{1}z+(1-t_{1})w$ and $c_{2}=t_{2}z+(1-t_{2})w$ this implies
$G_{j}(z)-G_{j}(w)=(\mathrm{Re}(G_{j}^{\prime}(c_{1}))+i\mathrm{Im}(G_{j}^{\prime}(c_{2})))^{T}(z-w).$
Similarly, we find $c_{3},c_{4}$ on the line segment from $w$ to $x$ with
$G(w)-G(x)=(\mathrm{Re}(G_{j}^{\prime}(c_{3}))+i\mathrm{Im}(G_{j}^{\prime}(c_{4})))(w-x).$
We have $z-w=i(\mathrm{Im}(z-x))$ and $w-x=\mathrm{Re}(z-x)$. Therefore,
$\mathrm{Re}(G_{j}(z)-G_{j}(x))$ $=-\mathrm{Im}(G_{j}^{\prime}(c_{2}))^{T}\,\mathrm{Im}(z-x)+\mathrm{Re}(G_{j}^{\prime}(c_{3}))^{T}\,\mathrm{Re}(z-x)$ (6)
$=\mathrm{Re}(a^{T}(z-x)),$
$\mathrm{Im}(G_{j}(z)-G_{j}(x))$ $=\mathrm{Re}(G_{j}^{\prime}(c_{1}))^{T}\,\mathrm{Im}(z-x)+\mathrm{Im}(G_{j}^{\prime}(c_{4}))^{T}\,\mathrm{Re}(z-x)$
$=\mathrm{Im}(b^{T}(z-x)),$
where $a=\mathrm{Re}(G_{j}^{\prime}(c_{3}))+i\mathrm{Im}(G_{j}^{\prime}(c_{2}))$ and $b=\mathrm{Re}(G_{j}^{\prime}(c_{1}))+i\mathrm{Im}(G_{j}^{\prime}(c_{4})).$ We get
$|G_{j}(z)-G_{j}(x)|\leq\sqrt{2}\,\max\{|a^{T}(z-x)|,|b^{T}(z-x)|\}\leq\sqrt{2}\,\max\{\|a\|_{1},\|b\|_{1}\}\,\|z-x\|_{\infty},$
where $\|\cdot\|_{1}$ is the $1$-norm (the dual of the $\infty$-norm). By construction, $c_{1},c_{2},c_{3},c_{4}$ are contained in $I$, because $I$ is convex. This implies that $a,b\in\mathrm{Re}(G_{j}^{\prime}(I))+i\mathrm{Im}(G_{j}^{\prime}(I)).$ It follows, using 5, that
$\max\{\|a\|_{1},\|b\|_{1}\}\leq\|\mathbf{1}_{n}-Y\cdot\square\mathrm{J}F(I)\|_{\infty},$
and so $\|G(z)-G(x)\|_{\infty}\leq\sqrt{2}\,\|\mathbf{1}_{n}-Y\cdot\square\mathrm{J}F(I)\|_{\infty}\,\|z-x\|_{\infty}.$
For the second part, we have to show that for all $I\in\mathbb{IC}^{n}$ we have $G(I)\subset K_{x,Y}(I)$. To show this we define the interval matrix $M:=(\mathbf{1}_{n}-Y\square\mathrm{J}F(I))\in\mathbb{IC}^{n\times n}$. By definition of $K_{x,Y}$ we have $G(x)+M(I-x)\subset K_{x,Y}(I)$. Thus, we have to show that $G(z)-G(x)\in M(I-x)$, since $z\in I$ is arbitrary. It follows from the equations in 6 and using 5 that we can find matrices $M_{1},M_{2}\in M$ such that $G(z)-G(x)=M_{1}\mathrm{Re}(z-x)+iM_{2}\mathrm{Im}(z-x)$. Decomposing the matrices into real and imaginary part we find
$G(z)-G(x)=\mathrm{Re}(M_{1})\mathrm{Re}(z-x)-\mathrm{Im}(M_{2})\mathrm{Im}(z-x)+i(\mathrm{Im}(M_{1})\mathrm{Re}(z-x)+\mathrm{Re}(M_{2})\mathrm{Im}(z-x)).$
Since $z-x\in I$ and by definition of the complex interval multiplication from (3) and the interval matrix-vector-multiplication (4) we see that $G(z)-G(x)\in M(I-x)$. This finishes the proof for the second part.
@@
$\lVert G_{Y}(z_{0})-G_{Y}(z_{1})\rVert_{\infty}\leq\sqrt{2}\lVert\mathbf{1}_{n}-Y\cdot\square\mathrm{J}F(I)\rVert_{\infty}\lVert z_{0}-z_{1}\rVert_{\infty}.$ (7)
By assumption $\sqrt{2}\lVert\mathbf{1}_{n}-Y\cdot\square\mathrm{J}F(I)\rVert_{\infty}$ is smaller than 1 so $G_{Y}$ is a contraction. Banach’s Fixed Point Theorem implies that $G_{Y}$ has a unique fixed point in $I$. Since $Y$ is invertible, this fixed point is the unique zero of $F$ in $I$. This shows the second part of Theorem 3.6.
Finally, let $(x_{i})_{i\geq 0}$ be the sequence defined in Proposition 3.10. By assumption, $x_{0}\in I$ and $x_{i+1}=G_{Y}(x_{i})$, $i\geq 0$. The second part of Lemma 6.1 together with an induction argument imply that $x_{i}\in I$ for all $i$. Let $x^{*}$ be the unique zero of $F$ in $I$. Then, 7 holds if we replace $z_{0}$ by $x^{*}$ and $z_{1}$ by $x_{i}$. We get $\lVert x^{*}-x_{i+1}\rVert_{\infty}<c\cdot\lVert x^{*}-x_{i}\rVert_{\infty}$ with a constant $c:=\sqrt{2}\lVert\mathbf{1}_{n}-Y\cdot\square\mathrm{J}F(I)\rVert_{\infty}<1$. This shows linear convergence and finishes the proof of Proposition 3.10. ∎
@@ 7 Certifying reality
Let $F:\mathbb{C}^{n}\rightarrow\mathbb{C}^{n}$ be a real square system of polynomials (or rational functions) and $I\in\mathbb{IC}^{n}$ a strong interval approximate zero of $F$. Then there exists $x\in I$ and $Y\in\mathbb{C}^{n\times n}$ satisfying $K_{x,Y}(I)\subset I$ and $\sqrt{2}\,\lVert\mathbf{1}_{n}-Y\square\mathrm{J}F(I)\rVert_{\infty}<1$. If additionally $\{\bar{z}\,|\,z\in K_{x,Y}(I)\}\subset I$, the associated zero of $I$ is real.